40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 6 10:04:58 2014, manasa, Update, General, Unexpected power shutdown 
    Reply  Mon Oct 6 10:23:43 2014, Steve, Update, VAC, Unexpected power shutdown 
       Reply  Tue Oct 7 16:55:16 2014, Steve, Update, VAC, Unexpected sweaty valves sweatyV5.jpgsweatyV5f.jpg
          Reply  Wed Oct 8 08:46:57 2014, Steve, Update, VAC, VAT valves actuator lubricant 
             Reply  Mon Oct 13 14:44:52 2014, Steve, Update, VAC, RGA scan pd78 -day 13 RGAscan13Oct2014.png
                Reply  Mon Dec 1 17:13:57 2014, Steve, Update, VAC, RGA scan pd78 -day 63 RGAscanD63.png
                   Reply  Tue Mar 31 09:03:32 2015, Steve, Update, VAC, RGA scan pd78 day 182 RGApd78d182.png
                      Reply  Wed Jul 1 10:01:25 2015, Steve, Update, VAC, RGA scan pd78 day 275 rga275d.png
                         Reply  Mon Oct 5 10:11:35 2015, Steve, Update, VAC, RGA scan pd78 day 386 RGA_background.png
       Reply  Tue Dec 23 14:33:24 2014, Steve, Update, VAC, Chiara gets UPS prepUP.png
    Reply  Mon Oct 6 10:28:18 2014, manasa, Update, General, Unexpected power shutdown 
       Reply  Mon Oct 6 11:09:52 2014, Jenne, Update, General, Unexpected power shutdown: slow computers 
       Reply  Mon Oct 6 17:04:51 2014, ericq, Update, General, Unexpected power shutdown 
          Reply  Mon Oct 6 17:36:17 2014, ericq, Update, General, Unexpected power shutdown 
          Reply  Mon Oct 6 18:15:12 2014, manasa, Update, General, Unexpected power shutdown: end green alignment 
    Reply  Tue Oct 7 16:17:15 2014, Steve, Update, General, reason for 8 sec power outage 
Message ID: 10573     Entry time: Mon Oct 6 18:15:12 2014     In reply to: 10571
Author: manasa 
Type: Update 
Category: General 
Subject: Unexpected power shutdown: end green alignment 

After Q brought back the IR, I went to check the green situation.

1. The end lasers had to be turned ON. 
2. The heaters for the doubler crystals had to be enabled. The heaters are at the set values.
3. The X arm PZTs for the steering mirrors had to be powered up (Set voltage 100V and current 6.7mA)
4. I aligned the green to the already IR-aligned arms.

Green PSL alignment has to be done after Q finishes his work on the MC WFS.

ELOG V3.1.3-